# Under-coverage in high-statistics counting experiments with finite MC samples

### v2

@@ 2.1.1 Alternate likelihood parametrization
which rotates the basis of the original parameters closer to the eigenvectors of the Hessian matrix. In fact, after applying the transformation in eq. (10), the linear correlation coefficient [^2] We remark that the condition number of the covariance matrix is invariant under a linear transformation such as eq. (10), which indeed offers a better metric to characterize the problem. between the transformed parameters becomes $0.71$ for the same value of $\epsilon$. However, we remark that the reduced correlation between parameters is not for free, given that the effective MC statistical power of the transformed templates $T_{ji}^{\prime}$ has also changed. The obvious solution would be to directly generate templates $T_{ji}^{\prime}$ for the transformed parameters, although this might not always be feasible for generic problems, for example because the MC generator is not flexible enough to allow for it.
When considering the alternate parametrization of eq. (10), we choose the POI (NP) to be $\mu_{1}^{\prime}$ ($\mu_{2}^{\prime}$) and change the definition of the alternate models accordingly, that is:
$\nu^{\pm}_{1i}=\nu_{1i}\cdot\left(1\pm 5\bar{\sigma}_{{\rm H}}^{\prime}\right),\;\;\;\nu^{\pm}_{2i}=\nu_{2i}\cdot\left(1\mp 5\bar{\sigma}_{{\rm H}}^{\prime}\right),$ (11)
where $\bar{\sigma}_{{\rm H}}^{\prime}$ is defined as the standard deviation of $\hat{\mu}_{1}^{\prime}$ for the nominal model in the limit of infinite-size MC samples. Given that $\mu_{1}$ and $\mu_{2}$ are highly correlated, it is easy to see that $\bar{\sigma}_{{\rm H}}^{\prime}\approx\bar{\sigma}_{{\rm H}}$, so for the sake of simplicity we will set $\bar{\sigma}_{{\rm H}}^{\prime}=\bar{\sigma}_{{\rm H}}$ in eq. (11). For this other parametrization, we have $\mu_{\rm t}=0$ for the nominal data and $\mu_{\rm t}=\pm 5\bar{\sigma}_{\rm H}$ for the alternate.
@@
The statistical uncertainty on $\mu_{2}^{\prime}$ is of ${\cal O}(N^{-1/2})$, so it will be highly constrained when $N$ is large. Hence, the uncertainty on ${\cal A}$, which is directly proportional to the uncertainty on the ratio $\mu^{\prime}_{1}/\mu_{2}^{\prime}$ as for eq. (12), is mostly driven by the limited precision on $\mu_{1}^{\prime}$. Measurements of asymmetries between two portions of a distribution occur frequently in Particle Physics. The determination of the $W$ boson mass at hadron colliders from the distribution of transverse momenta of the decaying leptons, already mentioned in section 1, is a relevant example of this kind. Indeed, this measurement can be interpreted, at least to a first approximation, as the determination of the asymmetry of the transverse momentum distribution around the Jacobian peak Rottoli:2023xdc .
2.1.2 Likelihood functions in the Gaussian approximation
@@
where the dots indicate terms that either do not depend on $f_{i}$ or are of higher-order in the expansion. Neglecting terms beyond the quadratic order amounts to approximate the joint Poisson density as a Gaussian multivariate with covariance matrix ${\bf V}={\rm diag}({\bf y})$. For the toy model, it has been verified that a joint Poisson density and its Gaussian approximation result in consistent point estimators and confidence intervals for all values of $N$ and $n$ considered in this study. For example, for $N=2\times 10^{6}$ and $n=200$, the number of events per bin is $y_{i}\sim 10^{4}$, and the second term in the expansion represents a correction of order $(y_{i}-\nu_{i})/y_{i}\sim\mathcal{O}(\sqrt{y_{i}})/y_{i}\sim 1\%$. The complete list of likelihood functions in the Gaussian approximation used in the study is presented in appendix A.
Another practical advantage of the Barlow-Beeston lite approach discussed in section 2.1 becomes evident in the Gaussian approximation. In order to see this, we first notice that the right-hand side of eq. (6) converges to a quadratic function of ${\bf f}$ in the limit $f_{i}\sim y_{i}\gg 1$, with Hessian matrix given by ${\bf V}^{-1}$. Profiling the $\boldsymbol{\beta}$ parameters in eq. (2.1) leads instead to a non-linear function of $(\theta,\mu)$, see e.g. eq. (72) in the appendix. However, the profiled likelihood function is still approximately quadratic in ${\bf f}$, with Hessian matrix given by $({\bf V}+{\bf V}_{\rm MC})^{-1}$, where ${\bf V}_{\rm MC}$ is a diagonal matrix with $i^{\rm th}$ diagonal-element equal to $f^{2}_{i}/m_{i}\approx y_{i}/k$, where $k$ is the relative proportion between data and simulation. This observation allows for an effective way to approximate the Barlow-Beeston lite likelihood as a quadratic function of its parameters, henceforth named “Gauss + MC stat.”, through a simple rescaling ${\bf V}\to{\bf V}\cdot(1+k^{-1})$. Numerical differences between minimizing the Barlow-Beeston lite likelihood or its quadratic approximation are found to be negligible in this context.
@@ 2.2 Confidence intervals in the asymptotic limit
Based on the properties of the maximum-likelihood estimators and of the profile-likelihood ratio in the asymptotic limit Kendall , we consider the following methods to set confidence intervals on $\mu$.
@@
    The Hessian uncertainty $\hat{\sigma}_{\rm H}$ is defined as the square root of the appropriate diagonal element of the inverse Hessian matrix of the NLL function at the minimum:
    ${\bf C}=\left(-\frac{\partial^{2}\ln L}{\partial^{2}(\mu,\boldsymbol{\theta})}\right)^{-1},$ (14)
    where $\boldsymbol{\theta}$ is a shorthand for the complete set of nuisance parameters. In the asymptotic limit, the maximum-likelihood estimators are known to be Gaussian-distributed around their true values Cowan , with covariance matrix given by the expectation value of ${\bf C}$, so that
@@
    The profile-likelihood ratio (PLR) is a function of $\mu$ defined as Kendall
@@
    where ${\hat{\boldsymbol{\theta}}}_{\mu}$ is the NP point that minimizes the NLL at a fixed value of $\mu$. In the asymptotic limit, the distribution of $t_{\mu}$ tends to a chi-square with one degree of freedom as for Wilks’ theorem Wilks , with deviations expected at the ${\cal O}(1/\sqrt{N})$ level, with $N$ related to the size of the data. If this high-order correction can be neglected, the interval
@@ 2.3 Numerical results
The minimization of the negative log-likelihood function is performed numerically using the MIGRAD minimizer implemented in Minuit2 JAMES1975343 , with PLR scans performed by MINOS and Hessian-based uncertainties computed by HESSE. Linear algebra operations are performed using the Eigen library eigenweb in double precision. The nominal configuration parameters for the toy model are reported in table 1. The effect of changing their values is studied in section 2.5. For simplicity, we have assumed $k_{1}=k_{2}=k$ [^3] For example, this would be appropriate when the signal and background can be defined as sub-processes of a same MC sample. The generalization to arbitrary $k_{j}$ is also straightforward and is not a critical aspect for this study.. Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu_{1}$ are obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments, and reported in table 2, while the results for the other choice $\mu=\mu^{\prime}_{1}$ are shown in table 3.
@@
in very good agreement with the numerical result shown in the first row of tables 2 and 3. Also, one can easily verify that the linear correlation between $\mu_{1}$ and $\mu_{2}$ differs from $-1$ by terms of ${\cal O}(\epsilon^{2})$. Differences between the Hessian uncertainty for the nominal and alternate models in table 2 are mostly due to the different number of expected events when $\mu_{\rm t}$ is changed.
$N\,\left[10^{6}\right]$ $\epsilon\,\left[\%\right]$ $n$ $k_{1}=k_{2}=k$
${\bf 2}$, $10$, $100$ ${\bf 3}$, $6$, $12$ $20$, $100$, ${\bf 200}$ ${\bf 1}$, $10$, $40$
@@
1) Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.095, $,\;0.095$ 0.095
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.106, $,\;0.106$ 0.106
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.083, $,\;0.083$ 0.083
2) Poisson Numeric Hessian 0 $0.521(5)$ 0.058, 0.0570.0580.0570.058,\;0.0570.058 , 0.057
$+5\bar{\sigma}_{{\rm H}}$ $0.160(4)$ 0.064, $,\;0.064$ 0.064
$-5\bar{\sigma}_{{\rm H}}$ $0.064(2)$ 0.050, $,\;0.050$ 0.050
3) Gauss Numeric Hessian 0 $0.520(5)$ 0.057, $,\;0.057$ 0.057
$+5\bar{\sigma}_{{\rm H}}$ $0.161(4)$ 0.064, $,\;0.064$ 0.064
$-5\bar{\sigma}_{{\rm H}}$ $0.063(2)$ 0.050, $,\;0.050$ 0.050
4) Gauss Analytic Hessian 0 $0.520(5)$ 0.057, $,\;0.057$ 0.057
$+5\bar{\sigma}_{{\rm H}}$ $0.161(4)$ 0.064, $,\;0.064$ 0.064
$-5\bar{\sigma}_{{\rm H}}$ $0.063(2)$ 0.050, $,\;0.050$ 0.050
5) Gauss + MC stat. Analytic Hessian 0 ${\bf 0.682}(5)$ 0.081, $,\;0.081$ 0.081
$+5\bar{\sigma}_{{\rm H}}$ $0.230(4)$ 0.086, 0.0850.0860.0850.086,\;0.0850.086 , 0.085
$-5\bar{\sigma}_{{\rm H}}$ $0.124(3)$ 0.076, $,\;0.076$ 0.076
6) Gauss + BB-lite Numeric Hessian 0 ${\bf 0.679}(5)$ 0.080, $,\;0.080$ 0.080
$+5\bar{\sigma}_{{\rm H}}$ $0.259(4)$ 0.094, $,\;0.094$ 0.094
$-5\bar{\sigma}_{{\rm H}}$ $0.097(3)$ 0.066, $,\;0.066$ 0.066
7) Gauss + BB Numeric Hessian 0 $0.439(5)$ 0.137, 0.1320.1370.1320.137,\;0.1320.137 , 0.132
$+5\bar{\sigma}_{{\rm H}}$ $0.449(5)$ 0.163, 0.1570.1630.1570.163,\;0.1570.163 , 0.157
$-5\bar{\sigma}_{{\rm H}}$ $0.454(5)$ 0.116, 0.1110.1160.1110.116,\;0.1110.116 , 0.111
8) Gauss + BB Numeric PLR 0 $0.454(5)$ 0.141, 0.1350.1410.1350.141,\;0.1350.141 , 0.135
$+5\bar{\sigma}_{{\rm H}}$ $0.461(5)$ 0.170, 0.1620.1700.1620.170,\;0.1620.170 , 0.162
$-5\bar{\sigma}_{{\rm H}}$ $0.469(5)$ 0.116, 0.1110.1160.1110.116,\;0.1110.116 , 0.111
9) Gauss + BB Numeric sPFC 0 $0.471(5)$ 0.160, 0.1460.1600.1460.160,\;0.1460.160 , 0.146
$+5\bar{\sigma}_{{\rm H}}$ $0.464(5)$ 0.184, 0.1690.1840.1690.184,\;0.1690.184 , 0.169
$-5\bar{\sigma}_{{\rm H}}$ $0.614(5)$ 0.166, 0.1550.1660.1550.166,\;0.1550.166 , 0.155
10) Gauss + BB Numeric PFC 0 $0.484(5)$ 0.172, 0.1570.1720.1570.172,\;0.1570.172 , 0.157
$+5\bar{\sigma}_{{\rm H}}$ $0.546(5)$ 0.211, 0.1880.2110.1880.211,\;0.1880.211 , 0.188
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.656}(5)$ 0.176, 0.1600.1760.1600.176,\;0.1600.176 , 0.160
11) Gauss + BB Numeric Barlett 0 $0.480(5)$ 0.168, 0.1550.1680.1550.168,\;0.1550.168 , 0.155
$+5\bar{\sigma}_{{\rm H}}$ $0.513(5)$ 0.205, 0.1840.2050.1840.205,\;0.1840.205 , 0.184
$-5\bar{\sigma}_{{\rm H}}$ $0.615(5)$ 0.166, 0.1520.1660.1520.166,\;0.1520.166 , 0.152
12) Gauss + BB Numeric CH 0 ${\bf 0.681}(5)$ 0.236, 0.2210.2360.2210.236,\;0.2210.236 , 0.221
$+5\bar{\sigma}_{{\rm H}}$ $0.776(4)$ 0.296, 0.2830.2960.2830.296,\;0.2830.296 , 0.283
$-5\bar{\sigma}_{{\rm H}}$ $0.847(4)$ 0.281, 0.2760.2810.2760.281,\;0.2760.281 , 0.276
13) Gauss + BB Numeric FC Cheat 0 ${\bf 0.664}(5)$ 0.229, 0.2070.2290.2070.229,\;0.2070.229 , 0.207
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.673}(5)$ 0.258, 0.2360.2580.2360.258,\;0.2360.258 , 0.236
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.680}(5)$ 0.190, 0.1710.1900.1710.190,\;0.1710.190 , 0.171
14) Gauss + BB Numeric Heuristic 0 ${\bf 0.678}(5)$ 0.228, 0.2210.2280.2210.228,\;0.2210.228 , 0.221
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.660}(5)$ 0.256, 0.2480.2560.2480.256,\;0.2480.256 , 0.248
$-5\bar{\sigma}_{{\rm H}}$ $0.720(4)$ 0.204, 0.1970.2040.1970.204,\;0.1970.204 , 0.197
@@
Table 2: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu_{1}$ for the toy model with $N=2\times 10^{6}$, $n=200$, $k=1$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.

### v3

@@ 2.2 Alternate likelihood parametrization
with $\mu^{\prime}_{1}$ ($\mu^{\prime}_{2}$) taken as POI (NP). Indeed, after applying the transformation in eq. (12), the linear correlation coefficient [^2] We remark that the condition number of the covariance matrix is invariant under a linear transformation such as eq. (12), which indeed offers a better metric to characterize the problem. between the transformed parameters changes to $0.71$ for the same value of $\epsilon$. However, we remark that the reduced correlation between parameters is not for free, given that the effective MC statistical power of the transformed templates
$T_{ji}^{\prime}=T_{1i}+(-1)^{j+1}T_{2i}$ (13)
has also changed. The obvious solution would be to directly generate templates $T_{ji}^{\prime}$ for the transformed parameters, although this might not always be feasible for generic problems, for example because the MC generator is not flexible enough to allow for it.
When considering the alternate parametrization of eq. (12), we thus set $\mu=\mu_{1}^{\prime}$ and $\theta=\mu_{2}^{\prime}$, and change the definition of the alternate data sets accordingly, that is:
$\nu^{\pm}_{1i}=\nu_{1i}\cdot\left(1\pm 5\bar{\sigma}_{{\rm H}}^{\prime}\right)\;\;\;{\rm and}\;\;\;\nu^{\pm}_{2i}=\nu_{2i}\cdot\left(1\mp 5\bar{\sigma}_{{\rm H}}^{\prime}\right),$ (14)
where $\bar{\sigma}_{{\rm H}}^{\prime}$ is defined as the standard deviation of $\hat{\mu}_{1}^{\prime}$ for the nominal model in the limit of infinite-size MC samples. Given that $\mu_{1}$ and $\mu_{2}$ are highly correlated, and that their variances are also similar for $\epsilon\ll 1$, it is easy to see that $\bar{\sigma}_{{\rm H}}^{\prime}\approx\bar{\sigma}_{{\rm H}}$, so for the sake of simplicity we will use $\bar{\sigma}_{{\rm H}}^{\prime}=\bar{\sigma}_{{\rm H}}$ in eq. (14). For this other parametrization, we have $\mu^{\prime}_{\rm t}=0$ for the nominal data and $\mu^{\prime}_{\rm t}=\pm 5\bar{\sigma}_{\rm H}$ for the alternate.
@@
The statistical uncertainty on $\mu_{2}^{\prime}$ is of ${\cal O}(N^{-1/2})$, so it will be highly constrained when $N$ is large. Hence, the uncertainty on $\cal A$}, which is directly proportional to the uncertainty on the ratio $\mu^{\prime}_{1}/\mu_{2}^{\prime}$ as for eq. (15), is mostly driven by the limited precision on $\mu_{1}^{\prime}$. Measurements of asymmetries between two portions of a distribution occur frequently in Particle Physics. The determination of the $W$ boson mass at hadron colliders from the distribution of transverse momenta of the decaying leptons, already mentioned in section 1, is a relevant example of this kind. Indeed, this measurement can be interpreted, at least to a first approximation, as the determination of the asymmetry of the transverse momentum distribution around the Jacobian peak [25].
2.3 Likelihood functions in the Gaussian approximation
@@
where the dots indicate terms that either do not depend on $f_{i}$ or are of higher-order in the expansion. Neglecting terms beyond the quadratic order amounts to approximate the joint Poisson density as a Gaussian multivariate with covariance matrix ${\bf V}={\rm diag}({\bf y})$. For the toy model, it has been verified that a joint Poisson density and its Gaussian approximation result in consistent point estimators and confidence intervals for all values of $N$ and $n$ considered in this study. For example, for $N=2\times 10^{6}$ and $n=200$, the number of events per bin is $y_{i}\sim 10^{4}$, and the second term in the expansion represents a correction of order $(y_{i}-\nu_{i})/y_{i}\sim\mathcal{O}(\sqrt{y_{i}})/y_{i}\sim 1\%$. The complete list of likelihood functions in the Gaussian approximation used in the study is presented in A.
Another practical advantage of the Barlow-Beeston lite approach discussed in section 2.1 becomes evident in the Gaussian approximation. To see this, we first notice that the right-hand side of eq. (7), which would be the correct NLL function in the case of infinite-size MC samples, converges to a quadratic function of $(\theta,\mu)$ in the limit $y_{i}\gg 1$, with error matrix given by $\bf V$}. Profiling the $\boldsymbol{\beta}$ parameters in eq. (2.1) introduces a nonlinearity in the parameters, which now appear also in the squared errors at the denominator of the profile likelihood function, see eq. (56) in the appendix. Formally, this function would be the same quadratic function as in the infinite-size MC case, provided that the error matrix is replaced by ${\bf V}+{\bf V}_{\rm MC}$, where ${\bf V}_{\rm MC}$ is a diagonal matrix with $i^{\rm th}$ diagonal element equal to $f^{2}_{i}/m_{i}$. This additional contribution to the error matrix depends on $(\mu,\theta)$, thus formally breaking the quadratic dependence. However, we notice that, for values of $(\mu,\theta)$ around the minimum, $f_{i}$ is constrained to take values of order $y_{i}$ by the penalty term $(y_{i}-f_{i})^{2}$ appearing at the numerator of the profile likelihood function, so that $f^{2}_{i}/m_{i}\approx y_{i}/k$ to a very good approximation, where $k\sim m_{i}/y_{i}$ represents the total statistical power of the MC simulation in the $i^{\rm th}$ bin relative to the data.
This observation allows us to conclude that the profiled likelihood for the Barlow-Beeston lite case, can be approximated, at least numerically, as a quadratic function of $\mu$ and $\theta$ with error matrix given by ${\bf V}+{\bf V}_{\rm MC}\approx{\bf V}\cdot(1+k^{-1})$. In the following, we will denote the latter function as “Gaus. + MC stat.”. As expected, numerical differences between minimizing the Barlow-Beeston lite likelihood or its Gaus. + MC stat. approximation are found to be negligible in the context of the toy study.
@@ 2.4 Confidence intervals in the asymptotic limit
Based on the properties of the maximum-likelihood estimators and of the profile-likelihood ratio in the asymptotic limit [28], we consider the following methods to set confidence intervals on $\mu$.
  • 1.
@@
    The Hessian uncertainty $\hat{\sigma}_{\rm H}$ is defined as the square root of the last diagonal element of the inverse Hessian matrix of the NLL function at the minimum:
    ${\bf C}=\left(-\frac{\partial^{2}\ln L}{\partial^{2}(\boldsymbol{\theta},\mu)}\right)^{-1},$ (17)
    where $\boldsymbol{\theta}$ is a shorthand for the complete set of nuisance parameters. In the asymptotic limit, the maximum-likelihood estimators are known to be Gaussian-distributed around their true values [10], with covariance matrix given by the expectation value of $\bf C$}, so that
@@
    The profile-likelihood ratio (PLR) is a function of $\mu$ defined as [28]
@@
    where $\hat{\mu}$ and $\hat{{\boldsymbol{\theta}}}$ are the maximum-likelihood estimators for $\mu$ and $\boldsymbol{\theta}$, respectively, while ${\hat{\boldsymbol{\theta}}}_{\mu}$ is the NP point that minimizes the NLL at a fixed value of $\mu$. In the asymptotic limit, the distribution of $t_{\mu}$ tends to a chi-square with one degree of freedom as for Wilks’ theorem [29], with deviations expected at the ${\cal O}(1/\sqrt{N})$ level, with $N$ related to the size of the data. If this high-order correction can be neglected, the interval
@@ 2.5 Numerical results
The minimization of the negative log-likelihood function is performed numerically using the MIGRAD minimizer implemented in Minuit2 [19], with PLR scans performed by MINOS and Hessian-based uncertainties computed by HESSE. Linear algebra operations are performed using the Eigen library [17] in double precision. The nominal configuration parameters for the toy model are reported in table 2. The effect of changing their values is studied in section 2.7. For simplicity, $k_{1}=k_{2}=k$ has been assumed [^3] For example, this would be appropriate when the signal and background can be defined as sub-processes of a same MC sample. The generalization to arbitrary $k_{j}$ is also straightforward and is not a critical aspect for this study.. Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu_{1}$ are obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments, and reported in table 3, while the results for the other choice $\mu=\mu^{\prime}_{1}$ are shown in table 4.
@@
in very good agreement with the numerical result shown in the first row of tables 3 and 4. Also, one can easily verify that the linear correlation between $\mu_{1}$ and $\mu_{2}$ differs from 1-1 by terms of ${\cal O}(\epsilon^{2})$. Differences between the Hessian uncertainty for the nominal and alternate models in table 3 are mostly due to the different number of expected events when $\mu_{\rm t}$ is changed.
$N\,\left[\times 10^{6}\right]$ $\epsilon\,\left[\times 0.01\right]$ $n$ $k_{1}=k_{2}=k$
$\bf 2$}, $10$, $100$ $\bf 3$}, 66, $12$ 2020, $100$, $\bf 200$} $\bf 1$}, $10$, $40$
@@
1) Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.095, 0.0950.095,\;0.095
$5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.106, 0.1060.106,\;0.106
$5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.083, 0.0830.083,\;0.083
2) Poisson Numeric Hessian 0 $0.521(5)$ 0.058, 0.0570.058,\;0.057
$5\bar{\sigma}_{{\rm H}}$ $0.160(4)$ 0.064, 0.0640.064,\;0.064
$5\bar{\sigma}_{{\rm H}}$ $0.064(2)$ 0.050, 0.0500.050,\;0.050
3) Gauss Numeric Hessian 0 $0.520(5)$ 0.057, 0.0570.057,\;0.057
$5\bar{\sigma}_{{\rm H}}$ $0.161(4)$ 0.064, 0.0640.064,\;0.064
$5\bar{\sigma}_{{\rm H}}$ $0.063(2)$ 0.050, 0.0500.050,\;0.050
4) Gauss Analytic Hessian 0 $0.520(5)$ 0.057, 0.0570.057,\;0.057
$5\bar{\sigma}_{{\rm H}}$ $0.161(4)$ 0.064, 0.0640.064,\;0.064
$5\bar{\sigma}_{{\rm H}}$ $0.063(2)$ 0.050, 0.0500.050,\;0.050
5) Gauss + MC stat. Analytic Hessian 0 ${\bf 0.682}(5)$ 0.081, 0.0810.081,\;0.081
$5\bar{\sigma}_{{\rm H}}$ $0.230(4)$ 0.086, 0.0850.086,\;0.085
$5\bar{\sigma}_{{\rm H}}$ $0.124(3)$ 0.076, 0.0760.076,\;0.076
6) Gauss + BB-lite Numeric Hessian 0 ${\bf 0.679}(5)$ 0.080, 0.0800.080,\;0.080
$5\bar{\sigma}_{{\rm H}}$ $0.259(4)$ 0.094, 0.0940.094,\;0.094
$5\bar{\sigma}_{{\rm H}}$ $0.097(3)$ 0.066, 0.0660.066,\;0.066
7) Gauss + BB Numeric Hessian 0 $0.439(5)$ 0.137, 0.1320.137,\;0.132
$5\bar{\sigma}_{{\rm H}}$ $0.449(5)$ 0.163, 0.1570.163,\;0.157
$5\bar{\sigma}_{{\rm H}}$ $0.454(5)$ 0.116, 0.1110.116,\;0.111
8) Gauss + BB Numeric PLR 0 $0.454(5)$ 0.141, 0.1350.141,\;0.135
$5\bar{\sigma}_{{\rm H}}$ $0.461(5)$ 0.170, 0.1620.170,\;0.162
$5\bar{\sigma}_{{\rm H}}$ $0.469(5)$ 0.116, 0.1110.116,\;0.111
9) Gauss + BB Numeric sPFC 0 $0.471(5)$ 0.160, 0.1460.160,\;0.146
$5\bar{\sigma}_{{\rm H}}$ $0.464(5)$ 0.184, 0.1690.184,\;0.169
$5\bar{\sigma}_{{\rm H}}$ $0.614(5)$ 0.166, 0.1550.166,\;0.155
10) Gauss + BB Numeric PFC 0 $0.484(5)$ 0.172, 0.1570.172,\;0.157
$5\bar{\sigma}_{{\rm H}}$ $0.546(5)$ 0.211, 0.1880.211,\;0.188
$5\bar{\sigma}_{{\rm H}}$ ${\bf 0.656}(5)$ 0.176, 0.1600.176,\;0.160
11) Gauss + BB Numeric Barlett 0 $0.480(5)$ 0.168, 0.1550.168,\;0.155
$5\bar{\sigma}_{{\rm H}}$ $0.513(5)$ 0.205, 0.1840.205,\;0.184
$5\bar{\sigma}_{{\rm H}}$ $0.615(5)$ 0.166, 0.1520.166,\;0.152
12) Gauss + BB Numeric CH 0 ${\bf 0.681}(5)$ 0.236, 0.2210.236,\;0.221
$5\bar{\sigma}_{{\rm H}}$ $0.776(4)$ 0.296, 0.2830.296,\;0.283
$5\bar{\sigma}_{{\rm H}}$ $0.847(4)$ 0.281, 0.2760.281,\;0.276
13) Gauss + BB Numeric FC Cheat 0 ${\bf 0.664}(5)$ 0.229, 0.2070.229,\;0.207
$5\bar{\sigma}_{{\rm H}}$ ${\bf 0.673}(5)$ 0.258, 0.2360.258,\;0.236
$5\bar{\sigma}_{{\rm H}}$ ${\bf 0.680}(5)$ 0.190, 0.1710.190,\;0.171
14) Gauss + BB Numeric Heuristic 0 ${\bf 0.678}(5)$ 0.228, 0.2210.228,\;0.221
$5\bar{\sigma}_{{\rm H}}$ ${\bf 0.660}(5)$ 0.256, 0.2480.256,\;0.248
$5\bar{\sigma}_{{\rm H}}$ $0.720(4)$ 0.204, 0.1970.204,\;0.197
@@
Table 3: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu_{1}$ for the toy model with $N=2\times 10^{6}$, $n=200$, $k=1$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative $\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.